OpenStudy (anonymous):

a car is traveling at 26mph. if its tires have a diameter of 25 inches , how fast are the car tires tunring???the answer in revolution per minutes...

hero (hero):

I would help you, but I'm about to go to bed. YAWN

hero (hero):

Somebody will help you

OpenStudy (anonymous):

hihiih thnkas... me too i go to bed now... i will come back here tomorrow:)... good nite:)

OpenStudy (anonymous):

answer is 364.33 revolution per min

OpenStudy (anonymous):

we need to deal with the circumference in thi sprob C=2 pi r = 2 pi 12 inch = 2 (3.14) 1 ft =6.28 ft now, speed is 26 miles per hour = 26 * 5280 ft per hour =26 * 5280 / 60 ft per min =2288 ft/min no of revolutions per minute = 2288 / 6.28 = 364.33

OpenStudy (anonymous):

linhphan if u find this answer correct and also u find it understandable, give medal

OpenStudy (dumbcow):

answer is 349.58 rev per min \[1 mile = (5280*12) inches\] 1 revolution is 25pi inches revolutions in 1 mile \[= \frac{5280*12}{25 \pi} = 806.72\] speed in miles per minute \[= \frac{26}{60} = 0.433\] revolutions per minute \[= 806.72 rev/mile*0.433 miles/\min = 349.58\]

OpenStudy (anonymous):

1 rev is not 25 pi inches dumbcow, check ur figures again correct answer is 364.33 and not 349.58

OpenStudy (dumbcow):

circumference = pi*diameter, i believe you set radius as 12 instead of 12.5 and that is why answers differ

OpenStudy (anonymous):

o correct dumbcow, i read diameter as 24 instead of 25 349.58 must be the right answer